Iron (III) oxide reacts with carbon to give iron and carbon monoxide.
Fe2O3 (s) + 3C(s) ….. 2Fe (s) + 3CO(s)

How many grams of Fe can be produced when 6.00 g of Fe2O3 reacts?
1
Expert's answer
2020-03-27T09:01:23-0400

Fe2O3

M = 159,69

2Fe

2M = 2*55.85 = 111.17

m (Fe) = 111.17*6.00/159,69 = 4,197 g - produced!